OpenStudy (legomyego180):

Simple question:

OpenStudy (legomyego180):

Does \[k^25^{-k^{3}}=\frac{ k^2 }{ 5^{-k^{3}} }\] or \[\frac{ k^2 }{ 5k^3 }\]

OpenStudy (legomyego180):

It's just a step im using in evaluating a improper integral and I forgot the exponent rules here

OpenStudy (mathmale):

It's the latter that is correct.

OpenStudy (legomyego180):

I know \[(x^a)^b=x^{a*b}\]

OpenStudy (legomyego180):

could you explain?

OpenStudy (mathmale):

Example:\[a ^{-1}=\frac{ 1 }{ a }\]

OpenStudy (mathmale):

\[k^25^{-k^{3}}=\frac{ k^2 }{ 5^{-k^{3}} }\]

OpenStudy (mathmale):

...is incorrect because you kept the negative sign in \[5^{-k^2}\]

OpenStudy (legomyego180):

whoops, didnt see that

OpenStudy (brydenwright):

take out the negative sign in your 1st option to make it correct

OpenStudy (mathmale):

If you re-write this as \[\frac{ 1 }{ 5^{k^2} }\]

OpenStudy (mathmale):

your original expression equals this result.

OpenStudy (mathmale):

What is the simplified version of -7y^(-2) ?

OpenStudy (legomyego180):

\(5k^3 \neq 5^{k^{3}}\) though

OpenStudy (mathmale):

The inequality symbol you used here is correct; the left side does not equal the right side.

OpenStudy (legomyego180):

\[\frac{ 7 }{ y^2 }\]

OpenStudy (mathmale):

The focus of this problem is what to do when you have a negative exponent and wish to simplify the expression by writing an equivalent expression with a positive exponent.

OpenStudy (legomyego180):

\(\color{blue}{\text{Originally Posted by}}\) @mathmale It's the latter that is correct. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (mathmale):

Note that there are 2 negative signs in -7y^(-2). Try again to simplify this. Goal: eliminate the negative exponent.

OpenStudy (legomyego180):

\[-7y^{-2} = \frac{ -7 }{ y^2 }\]

OpenStudy (mathmale):

Very good. You have eliminated the neg exponent properly, and you show recognition that the initial coefficient, -7, does not change.

OpenStudy (legomyego180):

Im still unsure how to solve the original problem though. I understand the rule here, but I still get \(\frac{ k^2 }{ 5^{k^{3}} }\) when I solve

zepdrix (zepdrix):

\[\large\rm \int\limits \frac{\color{royalblue}{k^2~dk}}{5^{\color{orangered}{k^3}}}\]Not seeing it? Should be a nice u-sub you can apply.

OpenStudy (legomyego180):

\[5^{-k^{3}}=\frac{ 1 }{ 5^{k^{3}} } = \frac{ 1 }{ 5^{3k} }\]

zepdrix (zepdrix):

No, that last step is no beuno.

zepdrix (zepdrix):

\[\large\rm a^{b^c}\ne a^{bc}\] \[\large\rm (a^b)^c=a^{bc}\]

OpenStudy (legomyego180):

\[(k^a)^b=k^{a*b} no?\]

OpenStudy (legomyego180):

ohh whoops

zepdrix (zepdrix):

The 3 is not being applied to the entire "thing", so you cannot apply that rule.

OpenStudy (legomyego180):

Still though, how does this simplify down to 5k^3?

zepdrix (zepdrix):

Honestly, you can just do u-sub from the start if you prefer.\[\large\rm \int\limits 5^{-k^3}k^2~dk\] \[\large\rm u=-k^3\]

zepdrix (zepdrix):

It doesn't. Mathmale must've misread your original statement.

OpenStudy (legomyego180):

OHHH ha, thats why I was confused

OpenStudy (legomyego180):

One of my classmates made the same mistake, so I was convinced it was right

zepdrix (zepdrix):

Ahh what a bozo :D

OpenStudy (legomyego180):

no worries I called him out on it lol